# SASG-DA: Sparse-Aware Semantic-Guided Diffusion Augmentation For Myoelectric Gesture Recognition

### v2

@@ Abstract
Surface electromyography (sEMG)-based gesture recognition plays a critical role in human–machine interaction (HMI), particularly for rehabilitation and prosthetic control. However, sEMG-based systems often suffer from the scarcity of informative training data, leading to overfitting and poor generalization in deep learning models. Data augmentation offers a promising approach to increasing the size and diversity of training data, where faithfulness and diversity are two critical factors to effectiveness. However, promoting untargeted diversity can result in redundant samples with limited utility. To address these challenges, we propose a novel diffusion-based data augmentation approach, Sparse-Aware Semantic-Guided Diffusion Augmentation (SASG-DA). To enhance generation faithfulness, we introduce the Semantic Representation Guidance (SRG) mechanism by leveraging fine-grained, task-aware semantic representations as generation conditions. To enable flexible and diverse sample generation, we propose a Gaussian Modeling Semantic Sampling (GMSS) strategy, which models the semantic representation distribution and allows stochastic sampling to produce both faithful and diverse samples. To enhance targeted diversity, we further introduce a Sparse-Aware Semantic Sampling (SASS) strategy to explicitly explore underrepresented regions, improving distribution coverage and sample utility. Extensive experiments on benchmark sEMG datasets, Ninapro DB2, DB4, and DB7, demonstrate that SASG-DA significantly outperforms existing augmentation methods. Overall, our proposed data augmentation approach effectively mitigates overfitting and improves recognition performance and generalization by offering both faithful and diverse samples.
@@ 1 Introduction
Gesture recognition serves as a fundamental technology for advancing human–machine interaction. Among various gesture recognition modalities, surface electromyography (sEMG)-based approaches have gained increasing attention due to their non-invasive nature, high temporal resolution, and ability to directly capture muscle activation signals associated with voluntary movement [kaifosh2025generic]. sEMG sensors can be conveniently worn on the skin surface, making them well-suited for continuous and user-friendly deployment in real-world scenarios, particularly in fields such as rehabilitation training, prosthetic control [xu2024chatemg], and wearable human–machine interfaces [sivakumar2024emg2qwerty].
@@
Data augmentation offers a promising approach to increasing the size and diversity of training data. Previous studies on sEMG-based gesture recognition have explored both single-sample augmentation [tsinganos2020data] and GAN-based augmentation [mendez2022emg, ao2024overcoming], which often suffer from limited diversity and unstable generation, respectively. More recently, diffusion-based augmentation has attracted significant interest for its ability to generate high-fidelity and diverse samples [dhariwal2021diffusion]. Nevertheless, despite its success in computer vision, only PatchEMG [xiong2024patchemg] has explored this direction in sEMG-based gesture recognition, focusing on generating high-quality data in a few-shot setting. In contrast, our work highlights two critical aspects of data augmentation: faithfulness and diversity [wang2024enhance, wang2025inversion, rahat2025data, trabucco2024effective]. Faithfulness ensures that generated samples remain consistent with the distribution of the original data, enabling the downstream model to learn from correct augmented data [wang2025inversion], while diversity enriches the original data distribution and helps the model learn robust and invariant features from diverse samples [wang2024enhance, wang2025inversion]. For instance, DiffMix [wang2024enhance] and AGA [rahat2025data] both emphasize introducing a diverse background while maintaining a faithful foreground, whereas DA-Fusion [trabucco2024effective] considers diversity in the textual prompt space. However, direct application of these approaches to sEMG signals remains limited due to the lack of explicit textual prompts and a clear foreground-background structure. Furthermore, promoting untargeted diversity can result in redundant samples that offer limited incremental value, an aspect that only a few studies [vu2025multi] have explored. Specifically, when diversity is introduced without considering the semantic coverage of the training set, the generated samples may cluster around already well-represented regions in the feature space, potentially reducing sample utility. Therefore, ensuring both faithfulness and targeted diversity helps expand the data distribution with informative samples, thereby improving the performance of sEMG-based gesture recognition systems (Fig. 1).
@@
To this end, we propose a novel diffusion-based augmentation method, Sparse-Aware Semantic-Guided Diffusion Augmentation (SASG-DA) for sEMG-based gesture recognition, which aims to generate both faithful and diverse augmented samples. To enhance generation faithfulness, we introduce a Semantic Representation Guidance (SRG) mechanism, which leverages fine-grained, class-specific representations as semantic conditions to guide the diffusion training process. These semantic representations capture rich contextual information, enabling the model to generate synthetic samples that are not only realistic but also well-aligned with the target classes. To enable flexible and diverse sample generation, we further propose a Gaussian Modeling Semantic Sampling (GMSS) mechanism for the diffusion inference process, which introduces diversity by stochastically sampling novel semantic conditions for generation. However, untargeted diversity may lead to redundant samples. To this end, we propose a Sparse-Aware Semantic Sampling (SASS) mechanism that actively targets sparse regions in the semantic space for conditional sampling, which explicitly expands the coverage of the training distribution. In this setting, diversity arises not only from the inherent stochasticity of the diffusion process but also from the stochastic variation introduced in the semantic condition space. Through the integrated strategy of Semantic Representation Guidance and Sparse-Aware Semantic Sampling, SASG-DA offers a principled augmentation method that enhances faithfulness and diversity, effectively mitigating overfitting and improving the performance and generalization of sEMG-based gesture recognition.
@@ 2 Related works
Surface electromyography (sEMG) signals reflect muscle activation patterns and have been widely used for gesture recognition, especially in human–computer interaction and rehabilitation applications [kaifosh2025generic]. Recent advances in deep learning have significantly improved sEMG-based gesture recognition by enabling end-to-end learning from raw multi-channel signals. Numerous studies have focused on designing tailored network architectures, such as Spatio-Temporal Cross Networks (STCNet) [yang2025stcnet], Kolmogorov-Arnold Network [al2024tcnn], and Transformer-based models [wang2024transformer], to enhance recognition accuracy. These models aim to capture the complex, non-stationary patterns in sEMG data and improve the performance and generalization of sEMG-based gesture recognition. In contrast, our work focuses on data augmentation as an orthogonal approach to enhance model performance and alleviate overfitting.
@@ 2.2 Data Augmentation for sEMG-based Gesture Recognition
Data augmentation is crucial in improving generalization and mitigating overfitting, especially in scenarios with limited training data. Existing sEMG augmentation strategies can be broadly categorized into single-sample DA and generative-based DA.
Single-sample sEMG augmentation methods apply various transformations to individual data instances while preserving labels, typically in the time or frequency domain [zhao2024dominant]. Prior work has systematically evaluated both traditional and novel data augmentation techniques for sEMG signals, such as magnitude warping and wavelet decomposition, demonstrating their effectiveness in improving classification accuracy [tsinganos2020data]. Although single-sample DA methods are simple and computationally efficient, they often offer limited diversity and insufficient variations for enhancing model generalization [cao2024survey].
Generative models provide a more principled approach to data augmentation by learning the underlying data distribution and synthesizing entirely new samples. Previous sEMG augmentation studies based on GANs [mendez2022emg, ao2024overcoming] and generative transformers (e.g., GPT) [bird2021synthetic] have demonstrated their effectiveness in mitigating data scarcity and improving robustness to signal variability and noise. Building upon these advancements, diffusion-based models have recently emerged as a compelling alternative, offering superior generation ability in terms of both sample diversity and fidelity [dhariwal2021diffusion]. However, only a few studies have explored their application to sEMG-based gesture recognition. PatchEMG [xiong2024patchemg] is the first to apply diffusion models to few-shot sEMG signal synthesis, achieving high-fidelity augmentation and improved gesture recognition accuracy. Nevertheless, it primarily focuses on synthesizing high-quality data from a minimal number of samples, rather than exploring broader augmentation strategies or enhancing sample diversity.
To the best of our knowledge, PatchEMG [xiong2024patchemg] and our work are currently the only studies that leverage diffusion models for sEMG data augmentation. While PatchEMG focuses on the few-shot setting, our work explores a broader augmentation approach, aiming to mitigate the overfitting and improve the performance of sEMG-based gesture recognition.
@@ 2.3 Diffusion Model Generation
Recent diffusion-based data augmentation methods primarily aim to improve the faithfulness and diversity of generated samples [wang2024enhance, wang2025inversion, rahat2025data, trabucco2024effective]. For instance, DiffuseMix [islam2024diffusemix] preserves key semantics by concatenating generated and original samples using a binary mask. Other methods emphasize enhancing diversity of generated samples [wang2024enhance, trabucco2024effective, rahat2025data]. Among them, DiffMix [wang2024enhance] interpolates inter-class images during inference to increase diversity, producing novel counterfactual samples such as land birds in maritime settings; DA-Fusion boosts diversity by inserting and fine-tuning new tokens in the text encoder to capture novel visual concepts [trabucco2024effective]. Beyond these efforts, only a few studies have explored the practical utility of augmented samples in downstream tasks [vu2025multi, li2025gendataagent]. As untargeted diversity may lead to redundant or typical samples, GenDataAgent [li2025gendataagent] incorporates downstream feedback to prioritize the generation of challenging samples. However, such feedback-driven approaches are often constrained by the slow inference speed of diffusion models. Moreover, most diffusion-based image augmentation methods are not readily applicable to sEMG signals due to their modality-specific design.
@@
A related but distinct objective is enhancing the diversity of diffusion generation, which has received comparatively less attention than improving generation quality [sadat2024cads]. Notable progress was made in CADS [sadat2024cads], where the authors showed that gradually annealing noise perturbations to conditional embeddings can significantly boost the diversity of generated samples. However, this approach is not intended for data augmentation.
@@
Minority sample generation targets low-density regions of the data distribution, since diffusion model samplers inherently favor high-density regions of the data manifold [sehwag2022generating]. Existing approaches typically address this limitation through gradient-based guidance, leveraging metrics such as the Hardness Score [sehwag2022generating] or Minority Score [ye2024don, um2024self, um2025minority] to generate minority samples. However, such gradient-guided methods often rely on additional external components and incur substantial computational overhead [um2024self].
@@
Distinct from prior works that focus solely on diverse or minority sample generation, our approach simultaneously emphasizes the faithfulness and diversity of generated data, which is crucial for effective data augmentation. Furthermore, our approach incorporates the Sparse-Aware Semantic Sampling (SASS) mechanism that explicitly expands the original data distribution to enhance augmentation utility, rather than focusing solely on rare samples in the target scenario. The latter approaches may generate data that deviates from the true distribution, which limits their usefulness for generalizable augmentation. Moreover, our SASS mechanism is designed to be computationally efficient.
@@
Figure 2: Overview of the proposed SASG-DA approach based on diffusion model. (b) SASG-DA enhances generation faithfulness by introducing semantic representations to guide diffusion training. (c) During inference, by (d) modeling semantic representation distribution, SASG-DA performs (e) Sparse-Aware Semantic Sampling to target the sparse regions in the semantic space and expand the data distribution. (f) The optimized sparse semantic representation are then used as condition inputs for diffusion reverse and generate synthetic sEMG signals for downstream model training.
@@ 3.1 Diffusion Models for Data Generation
Denoising Diffusion Probabilistic Models (DDPMs) approximate complex data distributions by modeling a Markovian forward diffusion process and learning its corresponding reverse denoising process [ho2020denoising]. In this work, we adapt the DDPM for sEMG-based gesture recognition, aiming to mitigate overfitting through data augmentation.
@@
The reverse process is a learned Markov chain that progressively denoises a sample from pure Gaussian noise to recover data resembling the original distribution. It is defined by $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are parameterized by a neural network $\theta$ [ho2020denoising].
Although the reverse process generates samples using the predicted means $\mu_{\theta}$, it is common in practice to express $\mu_{\theta}(x_{t},t)$ in terms of either the predicted clean data $\hat{x}_{0}(x_{t},t;\theta)$ or the predicted noise $\hat{\epsilon}(x_{t},t;\theta)$. Following the Diffusion-TS approach [yuan2024diffusion], our model is trained to predict an estimate $\hat{x}_{0}$ directly. This leads to a simplified training objective known as denoising score matching, formulated as:
@@
To further improve signal generative fidelity, we incorporate the same Fourier-based loss employed in Diffusion-TS [yuan2024diffusion]:
@@ 4.1 Overview
To mitigate the overfitting issue commonly encountered in sEMG-based gesture recognition, we propose a novel augmentation method, SASG-DA, based on diffusion models by offering faithful and diverse samples. To enhance generation faithfulness, SASG-DA conditions the diffusion model on fine-grained semantic representations extracted from the training data, ensuring that generated samples are both realistic and class-consistent. To enable flexible and diverse generation, we propose a Gaussian Modeling Semantic Sampling (GMSS) mechanism that approximates the semantic representation distribution based on a Gaussian assumption. To further explicitly expand the coverage of the training distribution and enhance data utility, we design a Sparse-Aware Semantic Sampling (SASS) mechanism, which encourages sampling from sparse or underrepresented regions of the modeled semantic space. The overall framework is illustrated in Fig. 2. Given a target class, our method first samples a semantic representation via SASS to guide the diffusion model in generating new gesture samples. These samples are then added to the training set to boost the performance of gesture recognition.

### v3

@@ Abstract
Surface electromyography (sEMG)-based gesture recognition plays a critical role in human–machine interaction (HMI), particularly for rehabilitation and prosthetic control. However, sEMG-based systems often suffer from the scarcity of informative training data, leading to overfitting and poor generalization in deep learning models. Data augmentation offers a promising approach to increasing the size and diversity of training data, where faithfulness and diversity are two critical factors to effectiveness. However, promoting untargeted diversity can result in redundant samples with limited utility. To address these challenges, we propose a novel diffusion-based data augmentation approach, Sparse-Aware Semantic-Guided Diffusion Augmentation (SASG-DA). To enhance generation faithfulness, we introduce the Semantic Representation Guidance (SRG) mechanism by leveraging fine-grained, task-aware semantic representations as generation conditions. To enable flexible and diverse sample generation, we propose a Gaussian Modeling Semantic Sampling (GMSS) strategy, which models the semantic representation distribution and allows stochastic sampling to produce both faithful and diverse samples. To enhance targeted diversity, we further introduce a Sparse-Aware Semantic Sampling strategy to explicitly explore underrepresented regions, improving distribution coverage and sample utility. Extensive experiments on benchmark sEMG datasets, Ninapro DB2, DB4, and DB7, demonstrate that SASG-DA significantly outperforms existing augmentation methods. Overall, our proposed data augmentation approach effectively mitigates overfitting and improves recognition performance and generalization by offering both faithful and diverse samples.
@@ 1 Introduction
Gesture recognition serves as a fundamental technology for advancing human–machine interaction (HMI). Among various gesture recognition modalities, surface electromyography (sEMG)-based approaches have gained increasing attention due to their non-invasive nature, high temporal resolution, and ability to directly capture muscle activation signals associated with voluntary movement [1]. sEMG sensors can be conveniently worn on the skin surface, making them well-suited for continuous and user-friendly deployment in real-world scenarios [2], particularly in fields such as rehabilitation training, prosthetic control, and wearable human–machine interfaces.
@@
Data augmentation offers a promising approach to increasing the size and diversity of training data. Previous studies on sEMG-based gesture recognition have explored both single-sample augmentation [3] and GAN-based augmentation [4], which often suffer from limited diversity and unstable generation, respectively. More recently, diffusion-based augmentation has attracted significant interest for its ability to generate faithful and diverse samples [5]. Nevertheless, despite its success in computer vision, only PatchEMG [6] has explored this direction in sEMG-based gesture recognition, focusing on generating high-quality data in a few-shot setting. In contrast, our work highlights two critical aspects of data augmentation: faithfulness and diversity [7, 8, 9]. Faithfulness ensures that generated samples remain semantically consistent with the distribution of the original data, enabling the downstream model to learn from correct augmented data [8], while diversity enriches the original data distribution and helps the model learn from robust and invariant features from diverse samples [7, 8]. For instance, DiffMix [7] emphasizes introducing a diverse background while maintaining a faithful foreground, whereas DA-Fusion [9] considers diversity in the textual prompt space. However, direct application of these approaches to sEMG signals remains limited due to the lack of explicit textual prompts and a clear foreground-background structure. Furthermore, promoting untargeted diversity can result in redundant samples that offer limited incremental value, an aspect that only a few studies [10] have explored. Specifically, when diversity is introduced without considering the semantic coverage of the training set, the generated samples may cluster around already well-represented regions in the feature space, potentially reducing sample utility. Therefore, ensuring both faithfulness and targeted diversity helps expand the data distribution with informative samples, thereby improving the performance of sEMG-based gesture recognition systems (Fig. 1).
@@
To this end, we propose a novel diffusion-based augmentation method, Sparse-Aware Semantic-Guided Diffusion Augmentation (SASG-DA) for sEMG-based gesture recognition, which aims to generate both faithful and diverse augmented samples. To enhance generation faithfulness, we introduce a Semantic Representation Guidance (SRG) mechanism, which leverages fine-grained, class-specific representations as semantic conditions to guide the diffusion training process. These semantic representations capture rich contextual information, enabling the model to generate synthetic samples that are not only realistic but also well-aligned with the target classes. Building upon SRG, we further propose a Gaussian Modeling Semantic Sampling (GMSS) mechanism for the diffusion inference process to enable flexible and diverse sample generation. GMSS introduces diversity by stochastically sampling novel semantic conditions for generation. However, untargeted diversity may lead to redundant samples. To this end, we further extend the GMSS by introducing a Sparse-Aware Semantic Sampling (SASS) mechanism that actively targets sparse regions in the semantic space for conditional sampling, which explicitly expands the coverage of the training distribution. In this setting, diversity arises not only from the inherent stochasticity of the diffusion process but also from the stochastic variation introduced in the semantic condition space. Through the integrated strategy of SRG and SASS, SASG-DA offers a principled augmentation method that enhances faithfulness and diversity, effectively mitigating overfitting and improving the performance and generalization of sEMG-based gesture recognition.
@@ 2 Related works
Surface electromyography (sEMG) signals reflect muscle activation patterns and have been widely used for gesture recognition, especially in human–computer interaction and rehabilitation applications [1]. Recent advances in deep learning have significantly improved sEMG-based gesture recognition by enabling end-to-end learning from raw signals. Numerous studies have focused on designing tailored network architectures, such as Spatio-Temporal Cross Networks (STCNet) [11], Kolmogorov-Arnold Network [12], and Transformer-based models [13], to enhance recognition accuracy. These models aim to capture the complex, non-stationary patterns of sEMG signals to enhance recognition performance and generalization. In contrast, our work focuses on data augmentation as an orthogonal approach to enhance model performance.
@@ 2.2 Data Augmentation for sEMG-based Gesture Recognition
Data augmentation is crucial in improving generalization and mitigating overfitting, particularly with limited training data. Existing sEMG augmentation strategies can be broadly categorized into single-sample DA and generative-based DA.
Single-sample sEMG augmentation methods [3] apply various transformations to individual data instances while preserving labels, typically in the time or frequency domain [14]. Although single-sample DA methods are simple and computationally efficient, they often offer limited diversity and insufficient variations for enhancing model generalization [15].
Generative models provide a more principled approach to data augmentation by learning the underlying data distribution and synthesizing entirely new samples. Previous sEMG augmentation studies based on GANs [4] have demonstrated their effectiveness in mitigating data scarcity and improving robustness. Building upon these advancements, diffusion-based models have recently emerged as a compelling alternative, offering superior generation ability in terms of both sample diversity and faithfulness [5]. However, only a few studies have explored their application to sEMG-based gesture recognition. PatchEMG [6] is the first to apply diffusion models to few-shot sEMG signal synthesis, focusing on generating high-quality data from limited samples and thereby improving gesture recognition accuracy. Specifically, it applies a patch-based training strategy to address the challenge of limited data for diffusion training, and during inference it generates complete-length signals and employs classifier-free guidance (CFG) [16] to ensure generation quality.
Given that EEG and EMG are both low-SNR, stochastic, and non-stationary biosignals [17], several representative EEG-oriented diffusion methods are conceptually relevant to sEMG data augmentation, including SSE-LDM [18], EEG-Enhancing [19], NTD [20] and DESAM [17]. Among them, DESAM [17] focuses on enhancing the faithfulness of generated samples through temporal mixup with real data. While these studies focus on improving the modeling and generation quality of diffusion models for EEG signals, they have not fully explored how to generate more effective augmented samples, relying mainly on the inherent diversity of diffusion models.
To the best of our knowledge, PatchEMG [6] and our work are among the first attempts that leverage diffusion models for sEMG data augmentation. While PatchEMG focuses on the few-shot setting, our work explores a broader augmentation approach to improve the downstream effectiveness of sEMG-based gesture recognition by emphasizing two critical factors: generation faithfulness and diversity.
@@ 2.3 Diffusion Model Generation
Recent diffusion-based augmentation methods primarily aim to improve the faithfulness and diversity of generated samples [7, 8, 9]. For instance, DiffuseMix [21] preserves key semantics by concatenating generated and original samples using a binary mask. Other methods emphasize enhancing diversity of generated samples [7, 9, 22]. Among them, DiffMix [7] interpolates inter-class images during inference to increase diversity, producing novel counterfactual samples such as land birds in maritime settings; DA-Fusion [9] boosts diversity by inserting and fine-tuning new tokens in the text encoder to capture novel visual concepts. Beyond these efforts, only a few studies have explored the practical utility of augmented samples in downstream tasks [10]. As untargeted diversity may lead to redundant or typical samples, GenDataAgent [10] incorporates downstream feedback to prioritize the generation of challenging samples. However, such feedback-driven approaches are often constrained by the slow inference speed of diffusion models. Moreover, most diffusion-based image augmentation methods are not readily applicable to sEMG signals due to their modality-specific design.
@@
A related but distinct objective is enhancing the diversity of diffusion generation, which has received comparatively less attention than improving generation quality [23]. Notable progress was made in CADS [23], which demonstrated that gradually annealing noise perturbations to conditions can significantly boost the diversity of generated samples. However, this approach is not intended for data augmentation.
@@
Minority sample generation targets low-density regions of the data distribution, since diffusion model samplers inherently favor high-density regions of the data manifold [24]. Existing approaches typically address this limitation through gradient-based guidance, leveraging metrics such as the Hardness Score [24] or Minority Score [25] to generate minority samples. However, such gradient-guided methods often rely on additional external components and incur substantial computational overhead [25].
@@
Distinct from prior works that focus solely on diverse or minority sample generation, our approach simultaneously emphasizes the faithfulness and diversity of generated data, which is crucial for effective data augmentation. Furthermore, our approach incorporates the Sparse-Aware Semantic Sampling (SASS) mechanism that explicitly expands the original data distribution to enhance augmentation utility, rather than focusing solely on rare samples in the target scenario. The latter approaches may generate data that deviates from the true distribution, which may limit their utility for augmentation. Moreover, our SASS mechanism introduces minor cost while requiring no additional components or gradient guidance.
@@
Figure 2: Overview of the proposed SASG-DA approach based on the diffusion model. (b) During diffusion training, SASG-DA enhances generation faithfulness by introducing semantic representation guidance. (c) During inference, SASG-DA models the semantic representation distribution with a Gaussian and performs Sparse-Aware Semantic Sampling to explicitly target sparse regions in the semantic space, thereby expanding the data distribution. (d) The optimized sparse semantic representations are then used as conditions for diffusion reverse and generate synthetic sEMG signals for downstream model training.
@@ 3.1 Diffusion Models for Data Generation
Denoising Diffusion Probabilistic Models (DDPMs) approximate complex data distributions by modeling a Markovian forward diffusion process and learning its corresponding reverse denoising process [26]. In this work, we adapt the DDPM for sEMG generation, aiming to mitigate the scarcity of informative training data.
@@
The reverse process is a learned Markov chain that progressively denoises a sample from pure Gaussian noise to recover data resembling the original distribution. It is defined by $p_{\theta}(x_{t-1}|x_{t})=\mathcal{N}(x_{t-1};\mu_{\theta}(x_{t},t),\Sigma_{\theta}(x_{t},t))$, where $\mu_{\theta}$ and $\Sigma_{\theta}$ are parameterized by a neural network $\theta$ [26].
In practice, it is common to express $\mu_{\theta}(x_{t},t)$ in terms of either the predicted clean data $\hat{x}_{0}(x_{t},t;\theta)$ or the predicted noise $\hat{\epsilon}(x_{t},t;\theta)$. Following the Diffusion-TS approach [27], our model is trained to predict an estimate $\hat{x}_{0}$ directly. This leads to a simplified training objective known as denoising score matching, formulated as:
@@
To further improve signal generative fidelity, we incorporate the same Fourier-based loss employed in Diffusion-TS [27] to better preserve the spectral structure of sEMG signals:
@@ 4.1 Overview
To mitigate the scarcity of informative training data commonly encountered in sEMG-based gesture recognition, we propose a novel augmentation method, SASG-DA, based on diffusion models by offering faithful and diverse samples. To enhance generation faithfulness, SASG-DA incorporates a Semantic Representation Guidance (SRG) mechanism, which conditions the diffusion model on fine-grained semantic representations extracted from the training data, ensuring that generated samples are both faithful and class-consistent. During diffusion inference, SRG is embodied through the proposed Gaussian Modeling Semantic Sampling (GMSS) and Sparse-Aware Semantic Sampling (SASS). To enable flexible and diverse generation, GMSS approximates the semantic representation distribution based on a Gaussian assumption. To further explicitly expand the coverage of the training distribution and enhance data utility, we extend the GMSS by designing a Sparse-Aware Semantic Sampling (SASS) mechanism, which encourages sampling from sparse or underrepresented regions of the modeled semantic space. The overall framework is illustrated in Fig. 2. Given a target class, our method first samples semantic representations via SASS to guide the diffusion model in generating new gesture samples. These samples are then added to the training set to boost the downstream performance of gesture recognition.